solve the systems by substitution x+2y=7
2x+2y=13

Answer by richwmiller(17219)

Solve the first equation for x and plug the result in for x in the second equation.
x+2y=7
x=7-2y
plug 7-2y in for x in the 2nd
2*(7-2y)+2y=13
now solve for y